OpenStudy (anonymous):

Find the derivative of f(x) = - 8 / x at x = 11.

OpenStudy (anonymous):

\[f(x)=-\frac{ 8 }{x }=-8 x ^{-1}\] \[f \prime(x)=-8(-1)x ^{-2}=\frac{ 8 }{ x^2 }\] \[f \prime(11)=?\]

OpenStudy (anonymous):

is f 8/x^2??

OpenStudy (anonymous):

\[f \prime(11)=\frac{ 8 }{ \left( 11 \right)^2 }=\frac{ 8 }{ 121 }\]

OpenStudy (anonymous):

okay that makes sense!! thanks so much:D

OpenStudy (anonymous):

yw